OpenStudy (anonymous):

Solve. x-5/3-2-x/5 = 17/15 a. Conditional b. Identity c. No Solution

OpenStudy (anonymous):

\[\frac{ x-5 }{ 3 }-\frac{ 2-x }{ 5 }=\frac{ 17 }{ 15 }\]

OpenStudy (anonymous):

it is not an identity for sure, because the expression on the left is not always equal to \(\frac{17}{5}\) it does have a solution do you need it?

OpenStudy (anonymous):

yes. I dont know how to start solving it

OpenStudy (anonymous):

clear the annoying fractions by multiplying both sides by the least common multiple of the denominators the least common multiple of 3,5,15 is 15 so start with \[15\left(\frac{ x-5 }{ 3 }-\frac{ 2-x }{ 5 }\right)=\frac{ 17 }{ 15 }\times 15\]

OpenStudy (anonymous):

use parentheses carefully \[5(x-5)-3(2-x)=17\] then solve as usual

OpenStudy (anonymous):

okay thanks :)